# VTCBench: Can Vision-Language Models Understand Long Context with Vision-Text Compression?

### v1

@@ Abstract
The computational and memory overheads associated with expanding the context window of LLMs severely limit their scalability. A noteworthy solution is vision-text compression (VTC), exemplified by frameworks like DeepSeek-OCR [51] and Glyph [9], which convert long texts into dense 2D visual representations, thereby achieving token compression ratios of 3x–20x. However, the impact of this high information density on the core long-context capabilities of vision-language models (VLMs) remains under-investigated. To address this gap, we introduce the first benchmark for VTC and systematically assess the performance of VLMs across three long-context understanding settings: VTC-Retrieval, which evaluates the model’s ability to retrieve and aggregate information; VTC-Reasoning, which requires models to infer latent associations to locate facts with minimal lexical overlap; and VTC-Memory, which measures comprehensive question answering within long-term dialogue memory. Furthermore, we establish the VTCBench-Wild to simulate diverse input scenarios. We comprehensively evaluate leading open-source and proprietary models on our benchmarks. The results indicate that, despite being able to decode textual information (e.g., OCR) well, most VLMs exhibit a surprisingly poor long-context understanding ability with VTC-compressed information, failing to capture long associations or dependencies in the context. This study provides a deep understanding of VTC and serves as a foundation for designing more efficient and scalable VLMs.
@@ 1 Introduction
In this work, we present the first systematic benchmark, VTCBench (Figure 1(b)), specifically designed for the VTC framework, rigorously quantifying the long-content comprehension capabilities of VLM models. Precisely, we focus on three critical tasks: (1) VTC-Retrieval requires models to retrieve, trace, and aggregate information (needles) placed at varying distances within a random text (haystack). (2) VTC-Reasoning tests the VLM’s capacity of associative reasoning [33] over a long context within the compressed visual space. The queried input has minimal literal matches with the context, thus assessing the model’s ability to reason associatively beyond mere lexical retrieval. (3) VTC-Memory assesses the VLM’s performance in very long-term dialogue memory [32, 53, 10, 62], evaluating the model’s resilience to temporal and structural degradation under vision-text compression. In addition, to fully capture the robustness against real-world visual variations, we introduce a challenging variant, VTCBench-Wild. These VTC-oriented long-context comprehension tasks establish a foundation to understand the effectiveness of VTC, offering crucial guidance for future design of highly efficient and capable next-generation long-context VLMs.
@@
  • Overall, existing VLM architectures exhibit weaker long-context comprehension abilities under the VTC framework compared to dedicated text-only LLMs. This indicates that existing VLM architectures have significant room for improvement when processing information compressed via VTC.
@@
  • In complex long-context understanding tasks, such as associative reasoning and long-term dialogue memory, the performance of existing VLMs under vision-text compression is notably poor, falling significantly below that of LLMs (e.g., Qwen3-8B [57]). This demonstrates that while VTC maintains satisfactory OCR performance, it seriously impedes core capabilities such as advanced associative reasoning and knowledge integration within long contexts.
  • To gain deeper insights into VLM failures, we conducted a qualitative error analysis that revealed recurring issues ranging from fundamental retrieval and visual-to-textual grounding to complex deficiencies in logical and associative reasoning.
  • Further experiments show that the VLM performance under VTC is critically dependent on both the font size and the spatial position of the information within the compressed image. We also establish VTCBench-Wild to simulate visual diversity encountered in real-world scenarios.
@@ 2.3 Long-Context Comprehension Benchmarks
The Needle-In-A-Haystack (NIAH) test is a classic method for evaluating long-context understanding capability of LLMs. The vanilla NIAH benchmark [25] evaluates simple retrieval robustness by embedding a specific fact (needle) within a long, irrelevant context (haystack). While useful for measuring basic context window functionality, it falls short of testing context understanding ability. Later, the NIAH task was enhanced by introducing multiple targets, adding distractor noise, and linking facts to enforce complex, or associative reasoning [20, 26, 43]. For example, NoLiMa [43] requires models to perform indirect inference, where the query and the required context do not share direct lexical overlap, forcing the model to rely on deeper semantic understanding and association maintenance across long distances. Beyond text-only benchmarks, NIAH in multimodal tasks has also been explored recently [48, 45]. Several benchmarks [32, 53, 10, 62] have been designed to evaluate long-term dialogue memory and understanding over very extended conversations.
@@ 5 VTCBench-Wild
Our ablation studies in Table 2 and the error analysis in Section 4.5 have demonstrated that the performance of VLMs is sensitive to various rendering parameters, such as font size, color, and line height. These findings indicate that evaluating models using a single, fixed rendering configuration may not fully capture their robustness to the visual diversity encountered in real-world documents. To bridge this gap and establish a more challenging evaluation standard, we introduce VTCBench-Wild, a new benchmark variant designed to simulate visually diverse “in-the-wild” scenarios. We construct VTCBench-Wild by sampling scenes from a rendering pool and questions from a question pool.
@@
Qwen3-8B (LLM baseline) [57] 99.38 62.25 54.17 73.55
@@
Experimental Results
We evaluated a diverse set of VLMs and an LLM baseline on VTCBench-Wild, with the comprehensive results presented in Table 7. A key finding is that the Qwen3-8B LLM baseline achieves the highest overall score (73.55%), driven by a near-perfect retrieval performance (99.38%). This suggests that for legible, albeit visually varied text, strong language understanding can be more effective than current VLM pipelines, which may introduce noise or errors during visual encoding.
Proprietary models lead the field, with Gemini-2.5-Pro (63.68%) and GPT-5 (61.23%) achieving the highest overall scores among all tested VLMs and outperforming their open-source counterparts. Among the two, Gemini-2.5-Pro exhibits stronger retrieval capabilities (87.38%) than GPT-5 (72.63%), though both fall short of the LLM baseline. Conversely, GPT-5 shows a slight edge in reasoning (62.88% vs. 58.75%). Both models, however, struggle significantly on the memory task, indicating that comprehending long, visually inconsistent conversational histories remains a major challenge.
The performance of open-source models reveals distinct architectural strengths and weaknesses. The Qwen3-VL series demonstrates exceptional retrieval scores (e.g., 97.50% for the 235B model), yet their reasoning performance is exceptionally poor (11.50% for 8B and 14.63% for 235B). This stark gap corroborates our error analysis in Section 4.5, highlighting a tendency to refuse associative reasoning. In contrast, its predecessor, Qwen2.5-VL-72B-Instruct, and Glyph show more balanced capabilities, achieving both high retrieval scores and substantially better reasoning performance. Notably, GLM-4.1V-9B-Thinking stands out by achieving the highest reasoning score among all VLMs, demonstrating superior robustness in multi-hop inference under diverse visual conditions. Conversely, models like InternVL3.5-8B and Gemma3-27B struggle across all tasks, with particularly low scores in reasoning and memory, underscoring their sensitivity to the visual diversity presented in VTCBench-Wild.
@@ 6 Conclusion
In this work, we conducted the first systematic investigation into the long-context understanding capabilities of VLMs under the vision-text compression (VTC) paradigm. We introduced comprehensive benchmarks VTCBench and VTCBench-Wild, which evaluate models on three complex tasks: information retrieval, associative reasoning, and long-term memory. Comprehensive experiments on a range of leading VLMs yielded valuable findings. While current models exhibit a respectable ability to perform simple text retrieval from compressed images, their performance is fragile, degrading with increased context length and compression ratios. More critically, we observed a near-total collapse in their ability to perform associative reasoning and maintain long-term memory, revealing a significant gap between visual perception and deep semantic comprehension. Furthermore, our analysis underscores that rendering parameters, particularly font size, are not a trivial implementation detail but a critical factor directly influencing model performance. High compression ratios achieved through smaller fonts severely impair even basic retrieval, confirming that VTC introduces unique perceptual challenges absent in text-only models.
@@ 7.2 Examples of VTCBench
VTC-Retrieval (NIAH) Lexical Matching, Multi-Hop Tracing, Aggregation, Question-Anwering (Dynamic query/key-value with types: word-word, word-number, uuid-number.) (essays) $\dots$ One of the special magic numbers for long-context is: 2026. $\dots$ One of the special magic numbers for distracting-information is: 2025. $\dots$ QA Variant. Q: What’s the special magic number for long-context? A: 2026. Completion Variant. Prompt: one of the special magic numbers for long-context is: Completion: 2026.
VTC-Reasoning (NIAH) Associative Reasoning, Question-Answering (Dynamic query/key-value with types: person-event/action.) (books) $\dots$ There was a vegan guest, named Katie. $\dots$ One-Hop Reasoning. Q: Which character cannot eat fish-based meals? A: Katie. Two-Hop Reasoning. Q: Which character cannot eat Brandade meals? A: Katie.

### v2

@@ Abstract
The computational and memory overheads associated with expanding the context window of LLMs severely limit their scalability. A noteworthy solution is vision-text compression (VTC), exemplified by frameworks like DeepSeek-OCR [55] and Glyph [10], which convert long texts into dense 2D visual representations, thereby achieving token compression ratios of 3x–20x. However, the impact of this high information density on the core long-context capabilities of vision-language models (VLMs) remains under-investigated. To address this gap, we introduce the first benchmark for VTC and systematically assess the performance of VLMs across three long-context understanding settings: VTC-Retrieval, which evaluates the model’s ability to retrieve and aggregate information; VTC-Reasoning, which requires models to infer latent associations to locate facts with minimal lexical overlap; and VTC-Memory, which measures comprehensive question answering within long-term dialogue memory. Furthermore, we establish the VTCBench-Wild to simulate diverse input scenarios. We comprehensively evaluate leading open-source and proprietary models on our benchmarks. The results indicate that, despite being able to decode textual information (e.g., OCR) well, most VLMs exhibit a surprisingly poor long-context understanding ability with VTC-processed information, failing to capture long associations or dependencies in the context. This study provides a deep understanding of VTC and serves as a foundation for designing more efficient and scalable VLMs.
@@ 1 Introduction
In this work, we present the first systematic benchmark, VTCBench (Figure 1(b)), specifically designed for the VTC framework, rigorously quantifying the long-content comprehension capabilities of VLM models. Precisely, we focus on three critical tasks: (1) VTC-Retrieval requires models to retrieve, trace, and aggregate information (needles) placed at varying distances within a random text (haystack). (2) VTC-Reasoning tests the VLM’s capacity of associative reasoning [37] over a long context within the compressed visual space. The queried input has minimal literal matches with the context, thus assessing the model’s ability to reason associatively beyond mere lexical retrieval. (3) VTC-Memory assesses the VLM’s performance in very long-term dialogue memory [36, 57, 11, 66], evaluating the model’s resilience to temporal and structural degradation under vision-text compression. In addition, to fully capture the robustness against real-world visual variations, we introduce a unified and lightweight variant, VTCBench-Wild. In summary, these VTC-oriented long-context comprehension benchmarks establish a foundation to understand vision-text compression, offering crucial guidance for future design of highly efficient and capable next-generation long-context VLMs.
@@
  • Overall, existing VLMs exhibit weaker long-context comprehension abilities under the VTC framework compared to dedicated text-only LLMs. This indicates that existing VLM architectures have significant room for improvement when processing information compressed via VTC.
@@
  • In complex long-context understanding tasks, such as associative reasoning and long-term dialogue memory, the performance of most VLMs utilizing VTC is notably poor, falling significantly below that of LLMs (e.g., Qwen3-8B [61]). This demonstrates that while VTC maintains satisfactory OCR performance, it seriously impedes core capabilities such as advanced associative reasoning and knowledge integration within long contexts.
  • Error analysis revealed that VLM failures under VTC range from visual-to-textual grounding to complex deficiencies in associative reasoning. Further experiments demonstrate that the performance is critically dependent on both the font size and the spatial layout of information within the compressed image.
  • Notably, Gemini-3-Pro [18] significantly outperforms some LLMs like Qwen3-8B [61] on VTCBench-Wild and nearly reaches parity with its own text-only baseline. This exceptional performance demonstrates that VTC is a highly viable and promising technical route for long-context understanding.
@@ 2.3 Long-Context Comprehension Benchmarks
The Needle-In-A-Haystack (NIAH) test is a classic method for evaluating the long-context understanding capability of LLMs. The vanilla NIAH benchmark [28] evaluates simple retrieval robustness by embedding a specific fact (needle) within a long, irrelevant context (haystack). While useful for measuring basic context window functionality, it falls short of testing context understanding ability. Later, the NIAH task was enhanced by introducing multiple targets, adding distractor noise, and linking facts to enforce complex, or associative reasoning [23, 29, 47]. For example, NoLiMa [47] requires models to perform indirect inference, where the query and the required context do not share direct lexical overlap, forcing the model to rely on deeper semantic understanding and association maintenance across long distances. Beyond text-only benchmarks, NIAH in multimodal tasks has also been explored recently [52, 49]. Several benchmarks [36, 57, 11, 66] have been designed to evaluate long-term dialogue memory and understanding over very extended conversations.
@@ 5 VTCBench-Wild
Our ablation studies in Table 2 and the error analysis in Section 4.5 have demonstrated that the performance of VLMs is sensitive to various rendering parameters, such as font size, color, and line height. These findings indicate that evaluating models using a single, fixed rendering configuration may not fully capture their robustness to the visual diversity encountered in real-world documents. To bridge this gap and establish a more challenging evaluation standard, we introduce VTCBench-Wild, a new benchmark variant designed to simulate visually diverse “in-the-wild” scenarios.
5.1 Construction of VTCBench-Wild
VTCBench-Wild is constructed by sampling scenes from a dynamic rendering pool and questions from a comprehensive question pool.
@@ 5.2 Evaluation Metrics
Qwen3-8B (Text Only) [61] 99.38 62.25 54.17 73.55
Gemini-3-Pro (Text Only) [18] 100.0 99.64 64.67 89.29
@@
Gemini-3-Pro [18] 99.88 97.63 58.00 87.64
Claude-Sonnet-4 [1] 68.25 8.75 36.83 38.05
Claude-Sonnet-4.5 [1] 76.00 10.88 42.83 43.27
Seed-1.5-Pro [20] 76.88 35.88 42.83 52.68
@@
5.3 Experimental Results and Analysis
We evaluated 17 VLMs and two text-only LLM baselines on VTCBench-Wild. The results, presented in Table 7, reveal a clear hierarchy of capabilities and highlight architectural differences in processing visually diverse, compressed textual information. We structure our analysis around three distinct groups, discussing both the challenges and the potential of the VTC paradigm:
  • LLM Baselines Illuminate Both the Gap and the Potential of VTC. The LLM baselines provide a crucial benchmark, revealing a persistent performance gap for the majority of VLMs. While text-only models like Qwen3-8B (73.55%) and Gemini-3-Pro (89.29%) set a high bar, most VLMs fail to match this level of performance, suggesting that the visual encoding stage often introduces an overhead or information loss. However, a critical exception to this trend is the Gemini-3-Pro VLM. With an overall score of 87.64%, it not only approaches the performance of its text-only counterpart but also decisively outperforms the strong Qwen3-8B LLM baseline. This standout result is highly significant: it serves as a compelling proof of concept for the VTC paradigm. It demonstrates that with a sufficiently advanced and well-integrated vision-language architecture, the challenges of visual compression can be overcome. Far from being an insurmountable barrier, this finding validates VTC as a viable, and potentially highly effective, pathway for scaling long-context processing.
  • Proprietary Models Lead, but Performance Varies Significantly. Among the VLMs, a clear stratification is evident. The VLM version of Gemini-3-Pro is the undisputed leader, achieving an overall score of 87.64% that nearly matches its text-only counterpart. This indicates a highly robust vision-language architecture capable of handling the visual variance in VTCBench-Wild with minimal performance degradation. Following at a considerable distance are Gemini-2.5-Pro (63.68%) and GPT-5 (61.23%). These models exhibit different strengths: Gemini-2.5-Pro shows stronger retrieval capabilities (87.38%), whereas GPT-5 has a slight edge in reasoning (62.88%). In contrast, other proprietary models like the Claude series show markedly lower performance, particularly on reasoning tasks, highlighting their sensitivity to the benchmark’s visual diversity.
  • Open-Source Models Reveal Distinct Architectural Profiles. The performance of open-source models uncovers diverse strengths and weaknesses. Glyph (57.27%) and Qwen2.5-VL-72B (57.36%) emerge as the most balanced and capable, achieving strong scores across retrieval and reasoning. A standout performer is GLM-4.1V-9B-Thinking, which, despite a modest overall score, achieves a remarkable 43.13% on the reasoning task. This is the highest among all open-source VLMs and is competitive even with leading proprietary models.
Overall, the results from VTCBench-Wild affirm that while many modern VLMs possess strong text perception abilities, this often does not translate to robust, deep contextual understanding. The Memory task proves to be a formidable challenge for nearly all models, with only Gemini-3-Pro scoring above 50%. The stark gap between retrieval and reasoning performance in models like Qwen3-VL highlights that true long-context comprehension requires more than just accurate OCR. VTCBench-Wild thus serves as an effective diagnostic tool, revealing the architectural resilience and specific failure points of VLMs in realistic scenarios.
@@ 6 Conclusion
In this work, we conducted the first systematic investigation into the long-context understanding capabilities of VLMs under the vision-text compression (VTC) paradigm. We introduced comprehensive benchmarks VTCBench and VTCBench-Wild, which evaluate models on three complex tasks: information retrieval, associative reasoning, and long-term memory. Comprehensive experiments on a range of leading VLMs yielded valuable findings. While current models exhibit a respectable ability to perform simple text retrieval from compressed images, their performance is fragile, degrading with increased context length and compression ratios. More critically, we observed a near-total collapse in their ability to perform associative reasoning and maintain long-term memory, revealing a significant gap between visual perception and deep semantic comprehension. Remarkably, with VTC, Gemini-3-Pro can outperform the LLM baseline (Qwen3-8B) on VTCBench-Wild. This exceptional performance validates VTC as a promising and viable paradigm for long-context comprehension. Furthermore, our analysis underscores that rendering parameters, particularly font size, are not a trivial implementation detail but a critical factor directly influencing model performance. High compression ratios achieved through smaller fonts severely impair even basic retrieval, confirming that VTC introduces unique perceptual challenges absent in text-only models.
@@ 7.2 Examples of VTCBench
VTC-Retrieval (NIAH) Lexical Matching, Multi-Hop Tracing, Aggregation (Dynamic query/key-value with types: word-word, word-number, uuid-number.) (essays) $\dots$ One of the special magic numbers for long-context is: 2026. $\dots$ One of the special magic numbers for distracting-information is: 2025. $\dots$ QA Variant. Q: What’s the special magic number for long-context? A: 2026. Completion Variant. Prompt: one of the special magic numbers for long-context is: Completion: 2026.
VTC-Reasoning (NIAH) Associative Reasoning, Question-Answering (Dynamic query/key-value with types: event/action-person.) (books) $\dots$ There was a vegan guest, named Katie. $\dots$ One-Hop Reasoning. Q: Which character cannot eat fish-based meals? A: Katie. Two-Hop Reasoning. Q: Which character cannot eat Brandade meals? A: Katie.